# GSR-Bench: A Benchmark for Grounded Spatial Reasoning Evaluation via Multimodal LLMs

### v1

@@ Abstract
The ability to understand and reason about spatial relationships between objects in images is an important component of visual reasoning. This skill rests on the ability to recognize and localize objects of interest and determine their spatial relation. Early vision and language models (VLMs) have been shown to struggle to recognize spatial relations. We extend the previously released What’sUp dataset Kamath et al. (2023) and propose a novel comprehensive evaluation for spatial relationship understanding that highlights the strengths and weaknesses of 27 different models. In addition to the VLMs evaluated in What’sUp, our extensive evaluation encompasses 3 classes of Multimodal LLMs (MLLMs) that vary in their parameter sizes (ranging from 7B to 110B), training/instruction-tuning methods, and visual resolution to benchmark their performances and scrutinize the scaling laws in this task.
GSR-Bench: A Benchmark for Grounded Spatial Reasoning Evaluation
via Multimodal LLMs
Navid Rajabi George Mason University nrajabi@gmu.edu                        Jana Košecká George Mason University kosecka@gmu.edu
@@ 1 Introduction
Earlier efforts for benchmarking vision and language models (VLMs) were developed for cross-modal and/or dual-encoder, end-to-end models, like LXMERT Tan and Bansal (2019), CLIP Radford et al. (2021), BLIP Li et al. (2022), with the focus on downstream tasks performances such as VQA Antol et al. (2015), GQA Hudson and Manning (2019), referring expressions Kazemzadeh et al. (2014), image-text matching or image/text retrieval.
@@
While spatial relations are often part of VQA datasets, the evaluation of spatial reasoning is often conflated with grounding referring expressions or objects and their attributes [^1] VQA example question may be: “Is there a woman to the left of the person that is wearing a wetsuit?”. To isolate these issues, authors in Kamath et al. (2023) introduced a new benchmark that focuses on spatial relationship understanding only. Using image-text matching evaluation methodology, they showed that contrastive models such as CLIP, BLIP, and their follow-up variants struggle to understand spatial relations with the best accuracy around $61$%.
Recent advances in generative large multi-modal models have shown remarkable visual knowledge and reasoning capabilities. We revisit the spatial relationship understanding in the context of MLLMs and extend the existing What’sUp benchmark Kamath et al. (2023) to include bounding box annotations and depth information. Compositional spatial relationship understanding requires successful recognition of objects and determining their locations. Furthermore, the knowledge of scene depth helps to disambiguate certain relationships (e.g., "in front of" or "behind"). The availability of this information can support a grounded understanding of spatial relations and will contribute to the fine-grained evaluation of large generative MLLMs, which lag behind their earlier counterparts. A few exceptions are multi-task multi-modal benchmarks like MMBench Liu et al. (2023d) and its related benchmarks that focus on evaluating several MLLMs for both visual recognition tasks and description generation. Given the simple structure of spatial clauses, we can study separately the ability of the model to ground the subject and object in the clause, and the effect and means of incorporating the depth information. The contributions of this work can be summarized as follows:
  • Extended What’sUp spatial relationship dataset with depth, masks, and bounding box annotations.
  • Design of different prompting strategies through structured prompting for the evaluation of grounding and spatial reasoning.
  • Comprehensive evaluation and comparison of 18 VLMs and 9 MLLMs, with various sizes, resolutions, pre-training/instruction-tuning, and prompting strategies.
@@
NegCLIP Yuksekgonul et al. (2022) 32.5 36.3 47.4 46.4 45.3 46.7 42.4
RoBERTaCLIP Kamath et al. (2023) 25.2 25.0 46.3 53.6 50.8 48.8 41.6
CoCa Yu et al. (2022) 2.1B 29.4 29.4 48.1 45.2 45.0 49.1 41.0
XVLM 4M Zeng et al. (2021) 216M 40.0 23.0 58.4 65.0 62.8 54.6 50.6
@@
Table 1: Template-based generation (TG) results using CircularEval. The first two sections come from What’sUp Kamath et al. (2023) results. The rest shows our LLaVA 1.5, 1.6, and InternVL-1.5 prompting results. Our best-performing is shown in bold, 2nd-best with underline, and What’sUp best-performing with italic underline.
@@ 2 GSR Benchmark
We extend carefully curated What’sUp dataset Kamath et al. (2023) that is comprised of Subset A containing pairs of objects in unambiguous spatial relations, being "on", "under", "left of" or "right of" a table, chair, or armchair, and Subset B containing an object "in front of", "behind", "left" or "right" of another object on a tabletop, and subsets of COCO-Spatial and GQA-Spatial with either one or two objects occurring, accompanied by spatial clauses like "on top of", "on the bottom of", "right of", or "left of". To study the grounding in this context, we annotate the dataset with bounding box coordinates and segmentation masks for all the objects mentioned in the captions and the depth maps for the images. We leverage GroundingDINO Liu et al. (2023c) as an open-vocabulary object detector, Segment Anything (SAM) Kirillov et al. (2023) for the object mask segmentation, and ZoeDepth Bhat et al. (2023) for monocular depth estimation. In the next section, we explain in detail how these additional annotations enable a more rigorous and grounded evaluation of spatial reasoning and its components [^2] All the code and data will be publicly available..
@@
3 GSR-Bench Experiments
Grounded spatial reasoning evaluation is typically done using image-text matching, binary VQA, or multiple-choice VQA. Further evaluations include subject and/or object grounding and localization; and exploring the effect of using depth information. In addition to 18 VLMs that have been evaluated in Kamath et al. (2023), we focus on the probing of open-source generative MLLMs like LLaVA and InternVL [^3] InternVL is the leading model in MMBench. using structured generation methodologies of Multiple choice (MC) and Template-based generation (TG). In MC prompting, captions for each image are represented as A, B, C, and D options for Subset A and Subset B, while A and B options for COCO-Spatial and GQA-Spatial Subsets. Then, the model is prompted to choose the correct letter as the final answer. In TG prompting, as shown in Figure 3, we append the correct format of the entire caption to the prompt, in which the spatial clause acts as the placeholder for the correct spatial relation option. In this way, we are able to leverage LLMs’ open-ended generation capability, handle the models’ verbosity by enforcing the correct answer structure, and overcome the biases observed in MC prompting simultaneously (See Figure 2).
Figure 3: TG sample prompt structure.
We ran each prompt with 4 different permutations so as to vary the position of the answer among the choices in MC and the list of options in TG prompting. An instance is considered correct if all four options are predicted correctly, known as CircularEval, introduced in MMBench Liu et al. (2023d). As opposed to the CircularEval, there exists VanillaEval, which only asks the model to choose the correct answer from a list of options once and has been shown to be prone to bias in recent studies. We first ran our experiments using MC prompting and observed a significant degree of bias among the models when the position of the answer varied among the choices of A, B, C, or D. This bias and sensitivity turned out to be even more detrimental in smaller models, while larger models like LLaVA-NeXT-Yi-34B and LLaMA-3-LLaVA-NeXT-8B showed significantly higher robustness (See Figure 4 in the Appendix for details). This phenomenon also corroborates the findings of multiple recent studies in LLMs Zheng et al. (2023); Pezeshkpour and Hruschka (2023); Wang et al. (2023); Xue et al. (2024); Wang et al. (2024). According to this observation, we opted for TG prompting, accompanied by the CircularEval methodology, inspired by Gemini 1.5 Pro Reid et al. (2024). See Table 1 for the TG prompting results, where rows in section 1 and 2 come from the What’sUp benchmark Kamath et al. (2023), section 3 refers to LLaVA-1.5 models Liu et al. (2023b), section 4 to the LLaVA-NeXT models Li et al. (2024); Liu et al. (2024), and section 5 to the InternVL-1.5 results Chen et al. (2024).
@@ 3 GSR-Bench Experiments
4 Discussion
According to Table 1 and 2, there is a positive correlation, even stronger in grounding, between scaling the LLM size & visual resolution, and the overall accuracy in both tasks. Conversely, there exist multiple exceptions, which are inevitable to concretely justify due to various intervening factors, such as (1) differences in training/fine-tuning & architectures and (2) release date and further instruction-tuning of the LLMs, like LLaMA-3-8B, which has the most-recent knowledge cut-off.
Grounding small objects, which refers to the Sub column in Subset A, seems challenging for all, and worst in smaller models, according to Table 2. We also observed a plateau in Table 1, especially in Qwen-1.5-110B, which is the largest ever released open-source MLLM at the moment. This could be a sign of saturation where the reasoning capability flattens out, although scaling still improves grounding, shown in Table 2.
@@ 5 Conclusions
Limitations
Small-scale Dataset:
Our split sizes remain the same as the What’sUp dataset in which Subset A has 412, Subset B has 408, COCO-Spatial-One has 2247, COCO-Spatial-Two has 440, GQA-Spatial-One has 1160, and GQA-Spatial-Two has 291 instances. Although this benchmark includes 4,958 image instances in total, each instance covering one or two objects, with various domain shifts in each 6 split, it is smaller than already existing benchmarks related to spatial reasoning, like Visual Genome Krishna et al. (2017), GQA Hudson and Manning (2019), VSR Liu et al. (2022), SpatialSense Yang et al. (2019), MMBench Liu et al. (2023d), etc. The reason is that this work aims to provide a carefully curated benchmark for spatial relationship understanding evaluation in a controlled setting to abstract away intervening factors that make the evaluations noisy.
Limited Spatial Prepositions:
@@
Lack of Robustness in MC Prompting:
@@
Intern-VL-1.5 Poor Grounding Observation:
@@
Depth Augmentation Nuances:
@@
Appendix A Appendix
The appendix is organized as follows:
  • Figure 4 demonstrates the biases of multiple-choice (MC) prompting.
  • Figures 5 - 11 depict the distributions of objects occurring in the captions.
  • Figure 12 shows sample failures in grounding small objects in Subset A.
@@ Appendix A Appendix
Figure 5: Subset A - subjects and objects distributions.
Figure 6: Subset B - subjects and objects distributions.
Figure 7: COCO Spatial One - objects distribution.
Figure 8: COCO Spatial Two - subjects distribution.
Figure 9: COCO Spatial Two - objects distribution.
Figure 10: GQA Spatial Two - subjects distribution.
Figure 11: GQA Spatial Two - objects distribution.

### v2

@@ Abstract
The ability to understand and reason about spatial relationships between objects in images is an important component of visual reasoning. This skill rests on recognizing and localizing objects of interest and determining their spatial relation. Early vision and language models (VLMs) have been shown to struggle to recognize spatial relations. We extend the previously released What’sUp [5] dataset and propose a novel comprehensive evaluation for spatial relationship understanding that highlights the strengths and weaknesses of 9 Multimodal LLMs (MLLMs), in comparison with the 18 VLMs tested in What’sUp dataset. Our experiments encompass three classes of MLLMs that vary in their parameter sizes (ranging from 7B to 110B), training/instruction-tuning methods, and visual resolution to benchmark their performances and scrutinize the scaling laws in this task.
@@ 1 Introduction
Earlier efforts for benchmarking vision and language models (VLMs) were developed for cross-modal and/or dual-encoder, end-to-end models, like LXMERT [23], CLIP [19], BLIP [10], with the focus on downstream tasks performances such as VQA [1], GQA [4], referring expressions [6], image-text matching or image/text retrieval. While spatial relations are often part of VQA datasets, the evaluation of spatial reasoning is often conflated with grounding referring expressions or objects and their attributes [^1] VQA example question may be: “Is there a woman to the left of the person that is wearing a wetsuit?”. To isolate these issues, authors in [5] introduced a new benchmark that focuses on spatial relationship understanding only. Using image-text matching evaluation methodology, they showed that contrastive models such as CLIP, BLIP, and their follow-up variants struggle to understand spatial relations with the best accuracy around $61$%. Recent advances in generative large multi-modal models have shown remarkable visual knowledge and reasoning capabilities. We revisit the spatial relationship understanding in the context of MLLMs and extend the existing What’sUp benchmark [5] to include bounding box annotations and depth information. Compositional spatial relationship understanding requires successful recognition of objects and determining their locations. Furthermore, the knowledge of scene depth helps to disambiguate certain relationships (e.g., "in front of" or "behind"). The availability of this information can support a grounded understanding of spatial relations and will contribute to the fine-grained evaluation of large generative MLLMs, which lag behind their earlier counterparts. A few exceptions are multi-task multi-modal benchmarks like MMBench [17] and its related benchmarks that focus on evaluating several MLLMs for both visual recognition tasks and description generation. Given the simple structure of spatial clauses, we can study separately the ability of the model to ground the subject and object in the clause, and the effect and means of incorporating the depth information. The contributions of this work can be summarized as follows: (1) Extended What’sUp spatial relationship dataset with depth, masks, and bounding box annotations, (2) Design of different prompting strategies through structured prompting for the evaluation of grounding and spatial reasoning, and (3) Comprehensive evaluation and comparison of 18 VLMs and 9 MLLMs, with various sizes, resolutions, pre-training/instruction-tuning, and prompting strategies.
@@
NegCLIP [30] 32.5 36.3 47.4 46.4 45.3 46.7 42.4
RoBERTaCLIP [5] 25.2 25.0 46.3 53.6 50.8 48.8 41.6
CoCa [29] 2.1B 29.4 29.4 48.1 45.2 45.0 49.1 41.0
XVLM 4M [31] 216M 40.0 23.0 58.4 65.0 62.8 54.6 50.6
@@
Table 1: Template-based generation (TG) results using CircularEval. The first two sections come from What’sUp [5] results. The rest shows our LLaVA 1.5, 1.6, and Intern-VL-1.5 prompting results. Our best-performing is bold, 2nd-best underlined, and What’sUp best-performing italic underlined.
@@ 2 GSR Benchmark
We extend carefully curated What’sUp dataset [5] that is comprised of Subset A containing pairs of objects in unambiguous spatial relations, being "on", "under", "left of" or "right of" a table, chair, or armchair, and Subset B containing an object "in front of", "behind", "left" or "right" of another object on a tabletop, and subsets of COCO-Spatial and GQA-Spatial with either one or two objects occurring, accompanied by spatial clauses like "on top of", "on the bottom of", "right of", or "left of". To study the grounding in this context, we annotate the dataset with bounding box coordinates and segmentation masks for all the objects mentioned in the captions and the depth maps for the images. We leverage GroundingDINO [16] as an open-vocabulary object detector, Segment Anything (SAM) [7] for the object mask segmentation, and ZoeDepth [2] for monocular depth estimation. In the next section, we explain in detail how these additional annotations enable a more rigorous and grounded evaluation of spatial reasoning and its components [^2] All the code and data will be publicly available..
3 GSR-Bench Experiments
Grounded spatial reasoning evaluation is typically done using image-text matching, binary VQA, or multiple-choice VQA. Further evaluations include subject and/or object grounding and localization; and exploring the effect of using depth information. In addition to 18 VLMs that have been evaluated in [5], we focus on the probing of open-source generative MLLMs like LLaVA and Intern-VL [^3] Intern-VL is the leading model in MMBench. using structured generation methodologies of Multiple choice (MC) and Template-based generation (TG). In MC prompting, captions for each image are represented as A, B, C, and D options for Subset A and Subset B, while A and B options for COCO-Spatial and GQA-Spatial Subsets. Then, the model is prompted to choose the correct letter as the final answer.
@@ 3 GSR-Bench Experiments
We ran each prompt with 4 different permutations so as to vary the position of the answer among the choices in MC and the list of options in TG prompting. An instance is considered correct if all four options are predicted correctly, known as CircularEval, introduced in MMBench [17]. As opposed to the CircularEval, there exists VanillaEval, which only asks the model to choose the correct answer from a list of options once and has been shown to be prone to bias in recent studies. We first ran our experiments using MC prompting and observed a significant degree of bias among the models when the position of the answer varied among the choices of A, B, C, or D. This bias and sensitivity turned out to be even more detrimental in smaller models, while larger models like LLaVA-NeXT-Yi-34B and LLaMA-3-LLaVA-NeXT-8B showed significantly higher robustness (See Figure 3 in the Appendix for details). This phenomenon also corroborates the findings of multiple recent studies in LLMs [32, 18, 24, 27, 25]. According to this observation, we opted for TG prompting, accompanied by the CircularEval methodology, inspired by Gemini 1.5 Pro [21]. See Table 1 for the TG prompting results, where rows in section 1 and 2 come from the What’sUp benchmark [5], section 3 refers to LLaVA-1.5 models [14], section 4 to the LLaVA-NeXT models [9, 15], and section 5 to the InternVL-1.5 results [3].
@@
Discussion.
According to Table 1 and 2, there is a positive correlation, even stronger in grounding, between scaling the LLM size & visual resolution, and the overall accuracy in both tasks. Conversely, there exist multiple exceptions, which are inevitable to concretely justify due to various intervening factors, such as (1) differences in training/fine-tuning & architectures and (2) release date and further instruction-tuning of the LLMs, like LLaMA-3-8B, which has the most-recent knowledge cut-off. Grounding small objects, which refers to the Sub column in Subset A, seems challenging for all, and worst in smaller models, according to Table 2. We also observed a plateau in Table 1, especially in Qwen-1.5-110B, which is the largest ever released open-source MLLM at the moment. This could be a sign of saturation in reasoning while grounding is still going up by scaling. Since the lowest grounding results have been achieved in localizing the subjects in Subset A, we also ran an additional Causal Effect analysis using linear regression to measure the causality between the grounding accuracy, as the independent variable, and the reasoning accuracy, as the dependent variable. As a result, we could reject the null hypothesis that "There is no causal effect between grounding and reasoning" due to the regression coefficient of $\beta$ = 6.025 and the $\rho$-value of 0.007, which is significantly lower than 0.05, shown in Figure 4.
@@ 4 Conclusions
Appendix A Appendix / supplemental material
A.1 Limitations
Small-scale Dataset.
Our split sizes remain the same as the What’sUp dataset in which Subset A has 412, Subset B has 408, COCO-Spatial-One has 2247, COCO-Spatial-Two has 440, GQA-Spatial-One has 1160, and GQA-Spatial-Two has 291 instances. Although this benchmark includes 4,958 image instances in total, each instance covering one or two objects, with various domain shifts in each 6 split, it is smaller than already existing benchmarks related to spatial reasoning, like Visual Genome [8], GQA [4], VSR [12], SpatialSense [28], MMBench [17], etc. The reason is that this work aims to provide a carefully curated benchmark for spatial relationship understanding evaluation in a controlled setting to abstract away intervening factors that make the evaluations noisy.
Limited Spatial Prepositions.
@@ A.1 Limitations
Lack of Robustness in MC Prompting.
@@
Intern-VL-1.5 Poor Grounding Observation.
@@
Depth Augmentation Nuances.
@@
A.2 Computational Resources
For LLaVA-1.5 (7B and 13B) and LLaVA-1.6/LLaVA-NeXT models up to the 13B versions, we used one 40 GB NVIDIA A100. For LLaVA-Yi-34B, and Intern-VL-Chat-1.5 26B, we used one 80 GB NVIDIA A100. For LLaVA-NeXT-Qwen-1.5 110B, we used four 80 GB NVIDIA A100 GPUs. In most cases, except the Intern-VL model, we used the Huggingface [26] library setup and checkpoints.
@@ A.2 Computational Resources
Figure 4: Causal effect analysis between grounding and reasoning accuracy, for the Subset A subjects that was the most difficult setting for the models for localization.